# LLM4SCREENLIT: Recommendations on Assessing the Performance of Large Language Models for Screening Literature in Systematic Reviews

### v1

@@ 4.2 Recommendations
Target Audience: Policymakers (Journals, Conferences, Guideline authors) Theme: Metrics and cost-sensitive evaluation (R1PM) Require reporting of Lost Evidence (Recall), MCC, and WMCC with declared FN:FP cost ratios, and discourage accuracy-centric or PABAK-focused reporting as primary evidence (origin: Section˜4.1). (R2PM) Mandate cost-sensitive evaluation narratives that explain trade-offs between efficiency and Lost Evidence, referencing WMCC or equivalent methods (origin: Section˜4.1). Theme: Reporting and transparency (R3PM) Require complete confusion matrices for all reported metrics to enable recomputation and meta-analytic synthesis (origin: Section˜4.1). (R4PM) Require disclosure of LLM output consistency and null-output rates, with explicit rules for handling unclassifiable or referred-back items in evaluation (origin: Sections˜4.1 and 2.4). (R5PM) Require open artifacts (prompts, seeds, code, data, and materials) and baseline comparators to support independent verification and fair comparison (origin: Section˜4.1 and (P2) in Section˜2.5). Theme: Study design and validity (R4PM) Require explicit leakage/contamination statements and temporal or provenance safeguards in retrospective or benchmark-based studies (origin: Section˜4.1, (P7) and (P8) in Section˜3.2). (R6PM) Require inclusion of non-LLM baselines for claims about efficiency or effectiveness, and disallow claims based on accuracy-only evidence (origin: Section˜4.1). Theme: Decision thresholds and governance (R7PM) Encourage pre-registration of acceptable Lost Evidence (minimum Recall) thresholds and escalation rules as part of protocol submissions, aligned to domain risk (origin: Section˜4.1).
@@ 4.3 Future research
This paper consolidates nine good practices (see Sections˜2.5 and 3.2) observed across the reviewed literature, demonstrates the pitfalls of accuracy-centric reporting under class imbalance, and proposes WMCC to integrate chance-correction with cost asymmetry, thereby turning fragmented results into operational guidance for SR screening, offering a coherent evaluation framework that supports more credible decisions. In spite of this, further research should investigate the effectiveness of combining multiple good practices identified by us (P1-P9) with ones identified by other researchers to keep the evaluation framework as robust as possible.
@@
Data will be available on request.
@@
The Matthews Correlation Coefficient (MMC) is a form of correlation coefficient calculated as:

### v2

@@ 4.2 Recommendations
Target Audience: Policymakers (Journals, Conferences, Guideline authors) Metrics and cost-sensitive evaluation   (R1PM) Require reporting of Lost Evidence (Recall), MCC, and WMCC with declared FN:FP cost ratios, and discourage accuracy-centric or PABAK-focused reporting as primary evidence (origin: Section˜4.1). (R2PM) Mandate cost-sensitive evaluation narratives that explain trade-offs between efficiency and Lost Evidence, referencing WMCC or equivalent methods (origin: Section˜4.1). Reporting and transparency   (R3PM) Require complete confusion matrices for all reported metrics to enable recomputation and meta-analytic synthesis (origin: Section˜4.1). (R4PM) Require disclosure of LLM output consistency and null-output rates, with explicit rules for handling unclassifiable or referred-back items in evaluation (origin: Sections˜4.1 and 2.4). (R5PM) Require open artifacts (prompts, seeds, code, data, and materials) to support independent verification and reproducibility (origin: Section˜4.1 and (P2) in Section˜2.5). Study design and validity   (R6PM) Require explicit leakage/contamination statements and temporal or provenance safeguards in retrospective or benchmark-based studies (origin: Section˜4.1, (P7) and (P8) in Section˜3.2). (R7PM) When the study design permits (e.g., SR updates, retrospective benchmarks with labelled datasets), encourage inclusion of non-LLM baselines for claims about efficiency or effectiveness; disallow claims based on accuracy-only evidence (origin: Section˜4.1). Decision thresholds and governance   (R8PM) Encourage pre-registration of acceptable Lost Evidence (minimum Recall) thresholds and escalation rules as part of protocol submissions, aligned to domain risk (origin: Section˜4.1).
4.3 Extending the Recommendations to Later SR Stages
Our recommendations were derived from evidence on title/abstract screening, but the underlying principles—class imbalance, asymmetric misclassification costs, and the need for chance-anchored metrics—are mathematical rather than stage-specific. Two of the papers in our review explicitly address later stages: Cao et al. [28] evaluate both abstract and full-text screening using prompt engineering strategies across multiple LLMs, and Khraisha et al. [9] report performance at three stages (title/abstract screening, full-text screening, and data extraction) using the same evaluation metrics. Drawing on these two papers and on the nature of each recommendation, Table˜6 summarises the applicability of R1–R10 across SR stages.
Table 6: Applicability of R1–R10 across SR stages. T = applies as stated; P = principle applies (the underlying principle carries over but the binary operationalisation requires adaptation—see footnotes); C = conditional (as for title/abstract screening).
R1 R2 R3 R4 R5 R6 R7 R8 R9 R10
[…]
aMCC and WMCC do not directly apply to structured information retrieval; the principles (chance-anchored, cost-asymmetric evaluation) transfer but require per-field or per-record metrics.
bRisk tolerance and escalation principles carry over, but thresholds must be defined per-field rather than per-study.
cThe analogue of a confusion matrix is a per-field accuracy table (correct, missing, and spurious extractions vs. a human-reference gold standard).
[…]
At the full-text screening stage, the task remains binary (include/exclude), so all recommendations transfer directly. Cao et al. [28] show that prompt engineering strategies (Framework CoT, ISO-ScreenPrompt) that differ from abstract-screening strategies can substantially improve full-text performance, but the evaluation metrics remain the same. Class imbalance is typically less extreme at full-text (both Cao et al. and Khraisha et al. [9] report higher inclusion ratios than for title/abstracts), however the asymmetric FN/FP costs remain.
At the data extraction stage, the task shifts from binary classification to structured information retrieval, so recommendations R1, R2, R3, R4, and R10 require not merely reinterpretation but a different evaluation design. Khraisha et al. [9] use the same metrics (Sensitivity, Specificity, Accuracy, Cohen’s kappa) for data extraction as for screening, treating each extraction field as a binary detection problem. While this is a reasonable first approximation, a richer evaluation would report per-field correctness, completeness, and spurious-extraction rates against a human-reference gold standard—an adaptation of R4 (confusion matrix) to the extraction context. Importantly, MCC generalises to multi-class problems via the $K{\times}K$ confusion matrix (see Gorodkin [47]), and WMCC can be naturally extended by the same logic: weighting each cell of the $K{\times}K$ confusion matrix by a corresponding entry in a $K{\times}K$ cost matrix (rather than a single scalar $w$) before applying Gorodkin’s formula. The principles that motivate these metrics—chance-anchored evaluation and asymmetric misclassification costs—thus carry over to multi-class extraction tasks, although the cost matrix must be elicited from domain stakeholders for each specific extraction context.
In the SE domain, data extraction tasks are frequently heterogeneous: mapping studies often involve multi-class categorisation of primary studies along facets such as research method, SE domain, contribution type, and venue type [22], SRs may require extracting quantitative data (effect sizes, sample sizes, and contextual variables), and quality assessment involves applying ordinal checklists rather than binary decisions. These tasks differ substantially from the binary per-field approach used by Khraisha et al. [9] and from biomedical extraction, where fields tend to be more structured (e.g., patient counts, drug doses, specific outcomes).
We emphasise that this applicability analysis is largely theoretical for later stages: the full-text screening row rests on two empirical sources (involving the same binary task), while the data extraction row rests on a single study plus the authors’ principled argument. Adapting the evaluation principles from this paper to these diverse SE extraction tasks—particularly the heterogeneous mapping, quantitative, and quality-assessment tasks described above—requires empirical validation that we flag as an open question for future work.
4.4 Guidance for Editors and Reviewers
The policymaker recommendations (R1PM–R8PM) specify what venues should require; this subsection offers practical guidance on how to enforce them.
Reviewer/Editor Checklist
The following checklist can be adopted by SE journals (such as IST) and conferences as part of their author guidelines or submission checklists for LLM-based SR screening studies. Authors should submit a completed copy of this checklist as part of the submission package, confirming which items are fulfilled and declaring any items that are not applicable with a brief justification. Fillable versions of this checklist in four formats (LaTeX, Word, Markdown, plain text), plus a PDF preview, are available in the replication package at https://doi.org/10.6084/m9.figshare.31356613.
LLM4SCREENLIT Reviewer/Editor Checklist $\square$ Complete confusion matrix (TP/FP/TN/FN) for every model $\times$ SR $\times$ prompt (R4, R3PM). $\square$ Lost Evidence ($1-\text{Recall}$) reported and discussed (R1, R1PM). $\square$ MCC reported unrescaled in [1,1][-1,1] (R1, R1PM). $\square$ WMCC with explicit weight $w$ and justification, even if $w=1$ (R1, R2, R2PM). $\square$ Null/invalid output rate, suspected cause, and handling rule specified (R6, R4PM). $\square$ Leakage/data-contamination statement with mitigation (R8, R6PM). $\square$ Replication package with prompts, seeds, code, and labelled data (R7, R5PM). $\square$ Deployment studies only: pre-specified Lost Evidence threshold with domain-stakeholder justification (R3) and escalation rule (R10, R8PM). $\square$ Benchmarking studies only: statement that deployment decisions are out of scope; R3/R10 are not required. $\square$ When the study aims to inform SR practice and training labels are available: non-LLM baseline comparison included (R9, R7PM).
Interpreting the Metrics
[…]
Table 7: Empirical performance categories for MCC and WMCC in SE LLM screening, derived from terciles of 183 evaluable configurations in the SESR-Eval reanalysis (9 LLMs $\times$ 24 SE secondary studies), following the approach of Kampenes et al. [48]. Medians are the median values within each tercile. These categories reflect the current state of LLM screening technology (2025 snapshot) and will shift as LLMs improve.
Small Medium Large
(lower tercile) (middle tercile) (upper tercile)
Range Median Range Median Range Median
MCC < 0.21{<}\,0.21 $0.09$ 0.210.210.380.38 $0.31$ > 0.38{>}\,0.38 $0.46$
WMCC (w=10w{=}10) < 0.19{<}\,0.19 $0.07$ 0.190.190.420.42 $0.30$ > 0.42{>}\,0.42 $0.56$
Approximately 8% of configurations had negative MCC values (worse than chance-level performance). Reviewers should note that WMCC << MCC is expected when $w>1$ (the FN penalty pulls the score down), and that WMCC >> MCC signals a classifier with unusually high Recall relative to its overall correlation. These categories should be treated as empirical reference points, not permanent thresholds (see Table˜7 caption).
Minimum Reporting Template
To illustrate what a minimally compliant results table should contain, Table˜8 shows a template that can be adapted by authors. Each row represents one model–prompt–SR combination; the columns cover the confusion-matrix counts and the metrics recommended in R1 and R2. Authors should provide one such table (or equivalent) for every evaluated configuration, together with the FN:FP cost ratio $w$ and any confidence intervals (R5) when sample-based validation was used.
Table 8: Minimum reporting template for an LLM-SR screening evaluation. Authors fill one row per model $\times$ prompt $\times$ SR combination. Italicised columns are deployment-only (R3, R5). The grey row shows a filled example from Table˜4.
Model Prompt SR TP FP FN TN Recall Lost Ev. MCC WMCC Null% CI
GPT-3.5-turbo-0613 PositiveX MobileMDE 47 50 8 187 0.855 14.5% 0.534 0.624 1.7%
Study type: Benchmarking / Deployment (delete improper). FN:FP cost ratio $w=$   (justification:  ).
[…]
Enforcement Actions
Concrete steps venues can take to operationalise the policymaker recommendations:
  1. 1.
    Strongly recommend (or even require) authors to submit a completed compliance declaration based on the checklist above, confirming that each applicable item is fulfilled and justifying any items marked as not applicable. Highlight this on the journal’s submission guidelines web page. This follows the practice widely adopted by medical journals, which require authors to submit completed PRISMA [45] checklists alongside their manuscripts. In the SE domain, SEGRESS [22] provides an analogous reporting checklist for secondary studies; our checklist extends this approach to LLM-based screening evaluations specifically.
  2. 2.
    Require complete confusion matrices as supplementary material at submission time (R3PM). Recommend that results tables follow the minimum reporting template (Table˜8) or an equivalent format covering all recommended columns.
  3. 3.
    Reject or require revision for submissions that lack any chance-anchored metric and report only Accuracy or PABAK as their primary classification performance measure (R1PM).
  4. 4.
    Require an explicit FN:FP cost-ratio declaration ($w$) in the method section; $w=1$ is acceptable only with justification (R2PM).
  5. 5.
    Require a leakage/contamination statement in retrospective or benchmark studies; flag papers using public benchmarks without one (R6PM). An example statement: “The SRs used for evaluation were published in [year range], [before/after] the training data cutoff of [LLM version]. We [did/did not] verify that the screened abstracts were not in the LLM’s training corpus. The risk of data contamination is [low/moderate/high] because [justification].”
  6. 6.
    Require a replication-package URL at submission (R5PM).
  7. 7.
    Adopt the checklist above as part of the venue’s author guidelines. Ready-to-use policy text (short and extended versions) for inclusion in author guidelines or calls for papers is available in the replication package at https://doi.org/10.6084/m9.figshare.31356613.
@@ 4.5 Future research
The LLM4SCREENLIT recommendations consolidate nine good practices (see Sections˜2.5 and 3.2) observed across the reviewed literature, demonstrate the pitfalls of accuracy-centric reporting under class imbalance, and propose WMCC to integrate chance-correction with cost asymmetry, thereby turning fragmented results into operational guidance for SR screening, offering a coherent evaluation framework that supports more credible decisions. In spite of this, further research should investigate the effectiveness of combining multiple good practices identified by us (P1-P9) with ones identified by other researchers to keep the evaluation framework as robust as possible. Also, retrospective studies of past SRs could include the use of additional human researchers performing the classification process to identify the actual performance of the current screening practice (i.e, two humans with disagreements resolved), as well as assessing the performance of using human-AI teams to generate classifications. See, for example, Cao et al. [28] who, although having data from several different completed SRs, based their study of prompt engineering method around samples from the different SR datasets and employed four human researchers to perform classifications on the samples to provide a rigorous assessment of both human and AI tool performance.
@@
Acknowledgements
We thank the anonymous reviewers for their thoughtful and constructive feedback, which led to significant improvements in the clarity, scope, and presentation of this paper.
@@ Acknowledgements
The replication package (extracted data, analysis scripts, and documentation) is available at https://doi.org/10.6084/m9.figshare.31356613.
@@
The Matthews Correlation Coefficient (MCC) is a form of correlation coefficient calculated as: